OpenStudy (anonymous):

Find the inverse of f(x) = [(2x+1)/(3x-1)]

OpenStudy (anonymous):

My work so far: y = [(2x+1)/(3x-1)] x = [(2y+1)/(3y-1)] x(3y-1) = (2y+1) I get lost here ^.

OpenStudy (anonymous):

isolate \(y\)

OpenStudy (anonymous):

\[3xy-x=2y+1\]\[3xy-2y=x+1\] ?

OpenStudy (anonymous):

Great, it looks clear to me now. So y = [(x+1)/(3x-2)] then? Pretty sure that's correct. Thanks again.

OpenStudy (anonymous):

very well :)